# Consistent Validation for Predictive Methods in Spatial Settings

### v2

@@ Abstract
Spatial prediction tasks are key to weather forecasting, studying air pollution impacts, and other scientific endeavors. Determining how much to trust predictions made by statistical or physical methods is essential for the credibility of scientific conclusions. Unfortunately, classical approaches for validation fail to handle mismatch between locations available for validation and (test) locations where we want to make predictions. This mismatch is often not an instance of covariate shift (as commonly formalized) because the validation and test locations are fixed (e.g., on a grid or at select points) rather than i.i.d. from two distributions. In the present work, we formalize a check on validation methods: that they become arbitrarily accurate as validation data becomes arbitrarily dense. We show that classical and covariate-shift methods can fail this check. We instead propose a method that builds from existing ideas in the covariate-shift literature, but adapts them to the validation data at hand. We prove that our proposal passes our check. And we demonstrate its advantages empirically on simulated and real data.
[…]
@@ 1 Introduction
In the spatial setting, predictive methods need not always arise from a statistical or machine learning approach built using training data. The predictive method is often a complex physical model provided by a third party (Remigio et al., 2022; Minnett, 2010; Gupta et al., 2018). Or it could combine physics and data-driven models (Banzon et al., 2016; Werner et al., 2019; Özkaynak et al., 2013).
In any of these cases, it is common to estimate the performance of a predictive method by using a set of validation data. More precisely, we are ultimately interested in predicting a response at what we call test sites; in the SST example above (Minnett, 2010), the test sites are points on a grid (often called a map), or in the air pollution example (Remigio et al., 2022), the test sites are 28 counties in the US Northeast. We can make predictions at the test sites, but we do not have access to direct observations of the responses there. We do have observed responses in the validation data, which we assume were not used in forming the predictive method being evaluated; in the SST example, scientists have SST observations taken by boats and buoys as validation data. So our aim is to estimate the average loss (i.e. risk) at the test sites using the validation data. We will see that many popular or natural approaches fail at this task.
One widely used approach, called the holdout, [^1] The name originally referred to “holding out” data for validation, with the remainder of available data going toward training. While we maintain the naming convention, we emphasize that in our setup there need not be any training data. Minnett (2010); Gupta et al. (2018); Duan et al. (2019), among many others, use this approach. estimates the test risk by taking the empirical average of the validation loss. When the validation and test data are independent and identically distributed (i.i.d.) from the same distribution, the holdout has a rigorous justification (Devroye, 1976; Langford, 2005). But in spatial problems, the validation and test sites need not be similarly dispersed, all data may be spatially correlated, and the test sites are often fixed rather than random; recall the grid or point prediction examples above. Indeed, Roberts et al. (2017) have observed problems with the holdout in practice. In the special case where the predictive method is data-driven, some authors (e.g. Telford and Birks, 2005) have suggested choosing holdout validation sites far from the training data sites. But we expect this proposal to still suffer from the problems just mentioned, and in fact simulation studies suggest it can misestimate test risk (Ploton et al., 2020; De Bruin et al., 2022).
@@
In what follows, we lay out a precise formulation of the prediction validation task in the spatial setting (Section 2). We formalize a desirable property for test-risk estimators: that, if arbitrarily dense validation data accrues in a region including the test points, the test-risk estimate should become arbitrarily accurate (Section 3). We prove that both the holdout estimator and an estimator advocated in the covariate-shift literature (Loog, 2012; Portier et al., 2023) fail to satisfy this spatial consistency property (Section 4). We propose to build on the $k$-nearest neighbor estimator (Loog, 2012). In particular, Loog (2012) and Portier et al. (2023) advocated fixing $k=1$ for covariate-shift problems. We instead derive an upper bound on the error of the general-$k$ estimator for estimating test risk (Section 5.1); crucially, our bound is conditional on the test and validation sites. We prove that choosing $k$ adaptively by optimizing our upper bound yields a spatially consistent estimator (Section 5.2). Unlike covariate-shift results (e.g. Portier et al., 2023), our results are directly applicable to problems where the test sites are most reasonably thought of as fixed. We illustrate the accuracy and practicality of our proposed method in simulated and real data analyses (Section 6), with tasks in both grid and point prediction. We discuss further related work in Appendix B.
2 Estimating Test Risk in a Spatial Problem
@@ 2.1 Test risk of a spatial predictive method
We assume we have access to a predictive method $h:(\mathcal{S},\mathcal{X})\to\mathcal{Y}$. We define $h^{\chi}:\mathcal{S}\to\mathcal{Y}$ to be the prediction made by $h$ at the location $S$: $h^{\chi}(S)=h(S,\chi(S))$. We suppose that practitioners would like to use $h$ to predict the response at a set of test sites where the response is unknown. We collect the $M^{\textup{{test}}}$ test data points, including true (but unobserved) responses, in $\smash{D^{\textup{{test}}}=(S^{\textup{{test}}}_{m},X^{\textup{{test}}}_{m},Y^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}}$.
To quantify quality of a predictive method, we need a loss. We assume the loss is bounded, as is often the case for practically-bounded responses; cf. temperature, pressure, or other physical quantities. [^2] Loss is also bounded for classification error or robust regression cases such as Tukey’s biweight loss.
@@
Due to practical considerations such as measurement error, the response at a test point is usefully modeled as random. To summarize loss over this randomness, it is standard to consider expected loss (a.k.a. risk) at the testing data. To define this expectation, we need to make assumptions about the data-generating process. It is typical in the non-spatial setting to assume responses are i.i.d. conditional on covariates. In the spatial setting, the i.i.d. assumption is inappropriate since it ignores spatial location. We instead assume that the response variable may be a function of the spatial location it is observed at, the covariates at that location, and i.i.d noise:
@@
Let $j=\text{test}$. Let $\chi:\mathcal{S}\to\mathcal{X}$ be a fixed function. For $1\leq m\leq M^{j}$, $X^{j}_{m}=\chi(S^{j}_{m})$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m},\epsilon_{m}^{j})$ with $f:\mathcal{S}\times\mathcal{X}\to\mathcal{Y}$ and $\epsilon_{m}^{j}\smash{\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}}P_{\epsilon}$ real-valued random variables.
Assumption 2.2 implies the response is i.i.d. given the spatial location. For example, Assumption 2.2 would cover the case where measurement errors on sensors are independent, but the locations of the sensors are not. A widely studied special case of Assumption 2.2 considers additive, homoskedastic noise: namely, $\mathcal{Y}\subset\mathbb{R}$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+\epsilon_{m}^{j}$. Assumption 2.2 is more general; for example, it allows the noise to be scaled by a continuous, deterministic function of $\mathcal{S}$: $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+g(S^{j}_{m})\epsilon_{m}^{j}$. Before defining risk, we first define the average loss of the predictive method at a particular location in space, $S$: $e_{h}(S):=\mathbb{E}[\ell(f(S,\chi(S),\epsilon),h^{\chi}(S))|S]$. Finally, we average over all spatial locations of interest, which we assume is a finite set.
@@ 2.2 Estimating test risk
$\smash{\hat{R}_{\textup{{Hold}}}(h)\coloneqq(1/N^{\textup{{val}}})\sum\nolimits_{n=1}^{N^{\textup{{val}}}}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n})).}$ (1)
For validation data to provide information about test risk, we need regularity assumptions. First, we make a standard assumption that validation data follows the same data-generating process as test data. And second, we assume a form of smoothness across the spatial locations.
@@
Assumption 2.2 remains true when we take $j=\text{val}$, with the same $f,\chi$, and $P_{\epsilon}$ as for $j=\text{test}$.
@@
Assumption 2.5 often arises naturally. For example, consider $\mathcal{Y}\subset[0,1]$, $\mathcal{S}=(\mathbb{R}^{d},\|\cdot\|_{2})$, squared loss, and homoskedastic and additive noise. Suppose $f(S,\chi(S))$ is $L_{Y}$-Lipschitz and $h^{\chi}$ is $L_{h}$-Lipschitz. Then Assumption 2.5 holds with $L=2(L_{Y}+L_{h})$; see Proposition D.1.
@@ 3 We Want Consistent Estimators
Once we have an estimator of test risk, it remains to check if that estimator performs well. We next formalize one natural check on performance: namely, estimators should become arbitrarily accurate if given validation data that is arbitrarily dense in the spatial domain. This check is analogous to traditional consistency in the i.i.d. data setting.
To that end, note that the fill distance is a measure of discrepancy between two sets of points, $\Psi_{1}$ and $\Psi_{2}$. [^3] The fill distance is not a distance in the mathematical sense since it is asymmetric and can equal 00 in cases when its two arguments are not exactly equal. It is the maximum distance from a point in $\Psi_{2}$ to the nearest point in $\Psi_{1}$.
@@
In the spatial statistics literature, infill asymptotics describes cases where data are gathered over a compact spatial domain in such a way that the fill distance of the data to its domain tends to 00 (Cressie, 2015, §5.8). We say an estimator is consistent for the test risk under infill asymptotics if – for any $Q^{\textup{{test}}}$, $\chi$, and $h$ satisfying our assumptions above – the estimator converges in probability to $R_{Q^{\textup{{test}}}}(h)$.
@@
That is, as validation data fills the spatial domain, the estimator should converge to the test risk – no matter the composition of test sites. Our assumption that fill distance tends to zero is generally weaker than an assumption that the validation sites are drawn i.i.d. from a distribution with Lebesgue density supported on the spatial domain. Reznikov and Saff (2015, Theorem 2.1) showed an implication relationship between these assumptions in a much more general setting, and Vacher et al. (2021, Lemma 12) discuss the special case for the unit cube. Next, we present a finite-sample version of this implication, with an advantage relative to past work that we keep track of all constants.
@@
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\smash{\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}}P$ for $1\leq n\leq N^{\textup{{val}}}$, and $P$ has Lebesgue density lower bounded by $c>0$ over $\smash{[0,1]^{d}}$. Let $B_{d}=\smash{\pi^{d/2}}/\Gamma(d/2+1)$ be the volume of the $d$-dimensional Euclidean unit ball. For any $\delta\in(0,1)$ there exists an $n_{0}$ such that for all $N^{\textup{{val}}}\geq n_{0}$ with probability at least $1-\delta$
@@
We prove Proposition 3.3 in Section D.2. The right side of this bound is $\smash{O((\log N^{\textup{{val}}}/N^{\textup{{val}}})^{1/d})}$, and so the fill distance converges to zero in probability under these assumptions.
@@
4 Current Estimators Exhibit Inconsistency
@@
While the holdout estimator of test risk is consistent for i.i.d. test and validation data (Devroye, 1976; Langford, 2005), we can construct examples showing Proposition 4.1 by observing that $\smash{\hat{R}_{\textup{{Hold}}}}$ has no dependence on the test task. So unless all test tasks have the same risk (which will be true only in unusually simplistic spatial settings), it cannot estimate them all consistently. The holdout estimator will generally exhibit non-trivial bias since it averages loss across the validation sites when we really care about loss at the test sites. See Section D.3.1 for a formal proof and also an example where the holdout converges to $\Delta$, the maximum possible error under the loss bound.
@@
Nearest Neighbor Estimator. Because of the assumed regularity in the error function (Assumption 2.5), it is natural to estimate the error at a test site using nearby validation points. Loog (2012) proposed risk estimators using $k$-nearest neighbors in the context of covariate shift. Both Loog (2012) and Portier et al. (2023) advocated for the use of 1-nearest neighbor (1NN) in the covariate-shift setting, with the latter providing theoretical justifications under standard covariate-shift assumptions. However, we show the 1NN estimator exhibits inconsistency in our spatial setting.
@@
We first review a general $k$-nearest neighbor estimator, which we revisit later. Define the $k$-nearest neighbor radius of a point $S\!\in\!\mathcal{S}$ as $\!\tau^{k}(S)\!:=\!\inf\{a\!\in\!\mathbb{R}\!:\!\!|S^{\textup{{val}}}_{1:N^{\textup{{val}}}}\!\cap\!B(S,a)|\!\geq\!k\},$ where $B(S,a)$ is the ball of radius $a$ centered at $S$. The $k$-nearest neighbor set [^4] We will state our results for the version of nearest neighbors where ties are resolved by including all equidistant points. However, our analysis holds for arbitrary tie-breaking methods. of a point $S\in\mathcal{S}$ is $A^{k}(S)\!:=\!\{1\!\leq n\!\leq\!N^{\textup{{val}}}\!:\!S^{\textup{{val}}}_{n}\in\!B(S,\!\tau^{k}(S))\}.$ As long as $1\leq k\leq N^{\textup{{val}}}$, $A^{k}(S)$ contains at least $k$ points: the $k$ nearest neighbors to $S$ in the validation set. It may be larger than $k$ if multiple points are equidistant from $S$.
@@
The $k$-nearest neighbor (kNN) test-risk estimator is defined by $\smash{\hat{R}_{\textup{{NN}},k}}(h):=\sum\nolimits_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))$, where $w^{\textup{{NN}},k}_{n}:=(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\mathrm{1}\{S^{\textup{{val}}}_{n}\!\in\!A^{k}(S^{\textup{{test}}}_{m})\}/|A^{k}(S^{\textup{{test}}}_{m})|.$
@@
For intuition, recall that—unlike in the covariate-shift setting—test points in the spatial setting are commonly fixed rather than arising i.i.d. from a distribution. Consider the simple case where $Q^{\textup{{test}}}=\delta_{S}$ for some $S\in\mathcal{S}$. Using $k=1$ leads us to estimate the error using a single validation point, which is inconsistent due to observation noise at the validation point. Where the problem with the holdout estimator was bias, the problem with 1NN is variance. In Section D.3.2 we prove Proposition 4.3 and show 1NN has large error when applied to classification point prediction tasks.
Inconsistency of Nearest Neighbors When $k$ Is a Function of the Number of Validation Points. In fact, we can prove a more general result: that any nearest-neighbor test-risk estimator where the number of neighbors depends (only) on the number of validation points is inconsistent under infill asymptotics, regardless of type of dependence.
@@
Let $(k_{n})_{n=1}^{\infty}$ be any sequence of natural numbers. Define the sequence of estimators $\smash{\hat{R}_{N^{\textup{{val}}}}}$ to be the nearest neighbor risk estimators using $N^{\textup{{val}}}$ validation points and $k_{N^{\textup{{val}}}}$ neighbors. Then there exists a data-generating process satisfying infill asymptotics, a test set containing a single point, a predictive method $h$ resulting in an error function satisfying the Lipschitz assumption, and an $\!\epsilon,\delta>\!0$ such that with probability at least $1\!-\!\delta$, $\forall N^{\textup{{val}}}$, $|\smash{\hat{R}_{N^{\textup{{val}}}}}(h)-R_{Q^{\textup{{test}}}}(h)|\geq\epsilon$.
See Section D.3.5 for a proof. There are two cases. (1) If the number of neighbors is bounded, the estimator suffers from non-vanishing variance as in the 1NN case. Or (2) the number is unbounded, so there exists a sequence of validation sites that accumulates slowly enough around each test site to lead to non-vanishing bias. Inconsistency of both 1NN and the holdout can be seen as corollaries of Proposition 4.4; for 1NN, choose: $\forall n,k_{n}=1$. For the holdout, choose: $\forall n,k_{n}=n$.

### v3

@@ Abstract
Spatial prediction tasks are key to weather forecasting, studying air pollution impacts, and other scientific endeavors. Determining how much to trust predictions made by statistical or physical methods is essential for the credibility of scientific conclusions. Unfortunately, classical approaches for validation fail to handle mismatch between locations available for validation and (test) locations where we want to make predictions. This mismatch is often not an instance of covariate shift (as commonly formalized) because the validation and test locations are fixed (e.g., on a grid or at select points) rather than i.i.d. from two distributions. In the present work, we formalize a check on validation methods: that they become arbitrarily accurate as validation data becomes arbitrarily dense. We show that classical and covariate-shift methods can fail this check. We propose a method that builds from existing ideas in the covariate-shift literature, but adapts them to the validation data at hand. We prove that our proposal passes our check. And we demonstrate its advantages empirically on simulated and real data.
@@ 1 INTRODUCTION
In the spatial setting, predictive methods need not always arise from a statistical or machine learning approach built using training data. The predictive method is often a complex physical model provided by a third party (Remigio et al., 2022; Minnett, 2010; Gupta et al., 2018). Or it could combine physics and data-driven models (Banzon et al., 2016; Werner et al., 2019; Özkaynak et al., 2013). In any of these cases, it is common to estimate the performance of a predictive method by using a set of validation data. More precisely, we are ultimately interested in predicting a response at what we call test sites; in the SST example above (Minnett, 2010), the test sites are points on a grid (often called a map), or in the air pollution example (Remigio et al., 2022), the test sites are 28 counties in the US Northeast. We can make predictions at the test sites, but we do not have access to direct observations of the responses there. We have observed responses in the validation data, which we assume were not used in forming the predictive method being evaluated; in the SST example, scientists have SST observations taken by boats and buoys as validation data. So our aim is to estimate the average loss (i.e. risk) at the test sites using the validation data. We show that many popular or natural approaches fail at this task.
One widely used approach, called the holdout, [^1] The name originally referred to “holding out” data for validation, with the remainder of available data going toward training. While we maintain the naming convention, we emphasize that in our setup there need not be any training data. Minnett (2010); Gupta et al. (2018); Duan et al. (2019), among many others, use this approach. estimates the test risk by taking the empirical average of the validation loss. When the validation and test data are independent and identically distributed (i.i.d.) from the same distribution, the holdout has a rigorous justification (Devroye, 1976; Langford, 2005). But in spatial problems, the validation and test sites need not be similarly dispersed, data may be spatially correlated, and the test sites are often fixed rather than random; recall the grid or point prediction examples above. Indeed, Roberts et al. (2017) observed problems with the holdout in practice. In the special case where the predictive method is data-driven, some authors (e.g. Telford and Birks, 2005) have suggested choosing holdout validation sites far from the training data sites. But this proposal still suffers from the problems just mentioned, and in fact simulation studies suggest it does not accurately measure prediction accuracy (Wadoux et al., 2021; De Bruin et al., 2022).
@@
In what follows, we start by laying out a precise formulation of the prediction validation task in the spatial setting (section 2). We formalize a desirable property for test-risk estimators: that, if arbitrarily dense validation data accrues in a region including the test points, the test-risk estimate should become arbitrarily accurate (section 3). We next prove that common estimators fail to satisfy this spatial consistency property (section 4); namely, we consider each of (1) the holdout estimator, (2) blocked spatial validation (Telford and Birks, 2005; Roberts et al., 2017), and (3) an estimator advocated in the covariate-shift literature (Loog, 2012; Portier et al., 2024). To construct a spatially consistent estimator, we build on the $k$-nearest neighbor estimator (Loog, 2012). Loog (2012) and Portier et al. (2024) advocated fixing $k=1$ for covariate-shift problems. We instead derive an upper bound on the error of the general-$k$ estimator for estimating test risk (section 5.1); crucially, our bound is conditional on the test and validation sites. We prove that choosing $k$ adaptively by optimizing our upper bound yields a spatially consistent estimator (section 5.2). Unlike covariate-shift results (e.g. Portier et al., 2024), our results are directly applicable to problems where the test sites are most reasonably thought of as fixed. We illustrate the accuracy and practicality of our proposed method in simulated and real data analyses (section 6), with tasks in both grid and point prediction. [^2] We provide an implementation of our method and the experiments in this paper at: https://github.com/DavidRBurt/Consistent-Spatial-Validation. We discuss further related work in appendix A.
2 SPATIAL TEST RISK
@@ 2.1 Test Risk of a Spatial Predictive Method
We assume we have access to a predictive method $h:\mathcal{S}\to\mathcal{Y}$. The predictive method might use the covariates. We allow for this dependence, but do not make it explicit in our notation. We suppose that practitioners would like to use $h$ to predict the response at a set of test sites where the response is unknown. We collect the $M^{\textup{{test}}}$ test data points, including true (but unobserved) responses, in $\smash{D^{\textup{{test}}}=(S^{\textup{{test}}}_{m},X^{\textup{{test}}}_{m},Y^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}}$.
To quantify quality of a predictive method, we need a loss. We assume the loss is bounded, as is common for responses naturally confined to a finite range; cf. temperature, pressure, or other physical quantities. [^3] Loss is also bounded for classification error or certain robust regression cases such as Tukey’s biweight loss.
@@
Due to practicalities such as measurement error, the response at a test point is usefully modeled as random. A standard summary of loss over this randomness is the expected loss (a.k.a. risk) at the test data. To define this expectation, we need to make assumptions about the data-generating process. It is typical in the non-spatial setting to assume responses are i.i.d. conditional on covariates. In the spatial setting, the i.i.d. assumption is inappropriate since it ignores spatial location. We instead assume that the response variable may be a function of the spatial location it is observed at, the covariates at that location, and i.i.d noise:
@@
Let $j\!\!=\!\textup{{test}}$, $\chi\!:\!\mathcal{S}\!\to\!\mathcal{X}$. For $1\!\leq\!m\!\leq\!M^{j}$, $X^{j}_{m}=\chi(S^{j}_{m})$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m},\epsilon_{m}^{j})$ with $f\!:\!\mathcal{S}\!\times\!\mathcal{X}\to\mathcal{Y}$ and $\epsilon_{m}^{j}\smash{\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}}P_{\epsilon}$ real-valued random variables.
Assumption 2.2 implies the response is i.i.d. given the spatial location. For example, assumption 2.2 covers the case where measurement errors on sensors are independent, but the locations of the sensors are not. A widely studied special case of assumption 2.2 considers additive, homoskedastic noise: namely, $\mathcal{Y}\subset\mathbb{R}$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+\epsilon_{m}^{j}$. Assumption 2.2 is more general; for example, it allows the noise to be scaled by a continuous, deterministic function of $\mathcal{S}$: $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+g(S^{j}_{m})\epsilon_{m}^{j}$.
Before defining risk, we first define the average loss of the predictive method at a particular location in space, $S$: $e_{h}(S):=\mathbb{E}[\ell(f(S,\chi(S),\epsilon),h(S))|S]$. Finally, we average over all spatial locations of interest, which we assume is a finite set.
@@
We write out the discrete distribution of test locations, $Q^{\textup{{test}}}$, to emphasize that test risk averages over the (fixed) test locations. We will find it useful for stating our upper bound on estimation error (theorem 5.1).
@@ 2.2 Estimating Test Risk
$\smash{\hat{R}_{\textup{{Hold}}}(h)\coloneqq\smash{\frac{1}{N^{\textup{{val}}}}}\sum\nolimits_{n=1}^{N^{\textup{{val}}}}\ell(Y^{\textup{{val}}}_{n},h(S^{\textup{{val}}}_{n})).}$ (1)
For validation data to provide information about test risk, we need to make an assumption about how validation data relates to test data. A particularly common (and often very reasonable) assumption is that validation and test data are i.i.d. However, as we argue above, the i.i.d. assumption is not appropriate in many spatial applications. To make progress, then, we need to make an alternative assumption relating validation and test data. Essentially we propose that, in spatial settings, smoothness in space is an alternative (and more intuitively plausible) regularity assumption.
More specifically, we first make a standard assumption that validation data follows the same data-generating process as test data, conditional on spatial locations. And second, we assume a form of smoothness in the average loss across the spatial locations.
@@
Assumption 2.2 remains true when we take $j=\textup{{val}}$, with the same $f,\chi$, and $P_{\epsilon}$ as for $j=\textup{{test}}$.
@@
Assumption 2.5 often arises naturally. For example, take $\mathcal{Y}\subset[0,1]$; $\mathcal{S}=(\mathbb{R}^{d},\|\cdot\|_{2})$; squared loss; and homoskedastic, additive noise. Suppose $f(S,\chi(S))$ is $L_{Y}$-Lipschitz and $h$ is $L_{h}$-Lipschitz. Then assumption 2.5 holds with $L=2(L_{Y}+L_{h})$; see proposition C.1.
@@ 3 WE WANT CONSISTENT ESTIMATORS
Given an estimator of test risk, we want to check if that estimator performs well. We next formalize one natural check on performance: namely, estimators should become arbitrarily accurate if given validation data that is arbitrarily dense in the spatial domain. This check is analogous to traditional consistency in the i.i.d. data setting.
To that end, the fill distance is a measure of discrepancy between two sets, $\Psi_{1}$ and $\Psi_{2}$. [^4] The fill distance is not a distance in the mathematical sense since it is asymmetric and can equal 00 in cases when its two arguments are not exactly equal. It is the maximum distance from a point in $\Psi_{2}$ to the nearest point in $\Psi_{1}$.
@@
In the spatial statistics literature, infill asymptotics describes cases where data are in a compact spatial domain so that the fill distance of the data to its domain tends to 00. We say an estimator is consistent for the test risk under infill asymptotics if—for any $Q^{\textup{{test}}}$, $\chi$, $f$ and $h$ satisfying our assumptions above—the estimator converges in probability to $R_{Q^{\textup{{test}}}}(h)$.
@@
That is, as validation data fills the spatial domain, the estimator should converge to the test risk—no matter the composition of test sites. Our assumption that fill distance tends to zero is generally weaker than an assumption that the validation sites are drawn i.i.d. from a distribution with Lebesgue density supported on the spatial domain. Reznikov and Saff (2015, Theorem 2.1) showed an implication relationship between these assumptions in a much more general setting, and Vacher et al. (2021, Lemma 12) discuss the special case for the unit cube. Next, we present a finite-sample version of this implication, with an advantage relative to past work that we keep track of all constants.
@@
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$, and $P$ has Lebesgue density lower bounded by $c>0$ over $\smash{[0,1]^{d}}$. Let $B_{d}=\pi^{d/2}/\Gamma(d/2+1)$ be the volume of the $d$-dimensional Euclidean unit ball. For any $\delta\in(0,1)$ there exists an $n_{0}$ such that for all $N^{\textup{{val}}}\geq n_{0}$ with probability at least $1-\delta$
@@
We prove proposition 3.3 in section C.2. The right side of this bound is $\smash{O((\log N^{\textup{{val}}}/N^{\textup{{val}}})^{1/d})}$, and so the fill distance converges to zero in probability under these assumptions.
@@
4 CURRENT ESTIMATORS ARE INCONSISTENT
@@
While the holdout estimator of test risk is consistent for i.i.d. test and validation data (Devroye, 1976; Langford, 2005), we can construct examples showing proposition 4.1 by observing that $\smash{\hat{R}_{\textup{{Hold}}}}$ has no dependence on the test task. So unless all test tasks have the same risk (which will be true only in unusually simplistic spatial settings), it cannot estimate them all consistently. The holdout estimator will generally exhibit non-trivial bias since it averages loss across the validation sites when we really care about loss at the test sites. See section C.3.1 for a formal proof and also an example where the holdout converges to $\Delta$, the maximum possible error for a bounded loss.
Inconsistency of Blocked Spatial Validation. We first define blocked spatial validation and then establish its spatial inconsistency. Imagine we have access to a set of data that may be used for training or validation. In blocked spatial validation, all of the data within a particular contiguous area (typically a block, or square) is reserved for validation, and the rest is used for training; see fig. 4 (left) in section C.3.4 for an illustration. Finally, the blocked spatial validation estimator of test risk is eq. 1; that is, this estimator is a special case of the holdout. While we focus on validation in the present work, blocked spatial validation often represents one fold within a cross-validation procedure (e.g. Milà et al., 2022; Linnenbrink et al., 2023; Le Rest et al., 2014; Wang et al., 2023); see section C.3.4 for a discussion. Prior work has shown via simulation that blocked spatial cross-validation can provide biased estimates of the test performance of a method (see Wadoux et al., 2021, Figure 1 and De Bruin et al., 2022, Figure 9). Since our counterexample in proposition 4.1 is not a blocked spatial validation estimator, it remains to show whether blocked spatial validation is spatially inconsistent.
Proposition 4.2 (Inconsistency of blocked spatial validation).
There exists a set of test points and a data-generating process satisfying infill asymptotics such that $\smash{\hat{R}_{\textup{{Hold}}}}$ applied in blocked spatial validation is not a consistent estimator of the test risk.
Spatial inconsistency arises for the same reason as in proposition 4.1: the validation locations may be systematically different from the test locations, and adding more validation locations need not resolve these differences. See section C.3.4 for a proof.
@@
Nearest Neighbor Estimator. Because of regularity in the error function (assumption 2.5), validation data are informative about the error at a nearby test site. Loog (2012) proposed risk estimators using $k$-nearest neighbors in the context of covariate shift. Both Loog (2012) and Portier et al. (2024) advocated for the use of 1-nearest neighbor (1NN) in the covariate-shift setting, with the latter providing theoretical justifications under standard covariate-shift assumptions. However, we show the 1NN estimator exhibits inconsistency in our spatial setting.
@@
We first review a general $k$-nearest neighbor estimator, which we revisit later. Define the $k$-nearest neighbor radius of a point $S\!\in\!\mathcal{S}$ as $\!\tau^{k}(S)\!:=\!\inf\{a\!\in\!\mathbb{R}\!:\!\!|S^{\textup{{val}}}_{1:N^{\textup{{val}}}}\!\cap\!B(S,a)|\!\geq\!k\},$ where $B(S,a)$ is the ball of radius $a$ centered at $S$. The $k$-nearest neighbor set [^5] We state our results for nearest neighbors with ties resolved by including all equidistant points. However, our analysis holds for arbitrary tie-breaking methods. of a point $S\in\mathcal{S}$ is $A^{k}(S)\!:=\!\{1\!\leq n\!\leq\!N^{\textup{{val}}}\!:\!S^{\textup{{val}}}_{n}\in\!B(S,\!\tau^{k}(S))\}.$ As long as $1\leq k\leq N^{\textup{{val}}}$, $A^{k}(S)$ contains at least $k$ points: the $k$ nearest neighbors to $S$ in the validation set. It may be larger than $k$ if multiple points are equidistant from $S$.
@@
The $k$-nearest neighbor (kNN) test-risk estimator is $\smash{\hat{R}_{\textup{{NN}},k}}(h)\!:=\!\sum\nolimits_{n=1}^{N^{\textup{{val}}}}\!w^{\textup{{NN}},k}_{n}\ell(Y^{\textup{{val}}}_{n},h(S^{\textup{{val}}}_{n}))$, with $w^{\textup{{NN}},k}_{n}\!:=\frac{1}{M^{\textup{{test}}}}\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\mathrm{1}\{S^{\textup{{val}}}_{n}\!\in\!A^{k}(S^{\textup{{test}}}_{m})\}/|A^{k}(S^{\textup{{test}}}_{m})|$.
@@
For intuition, recall that—unlike in the covariate-shift setting—test points in the spatial setting are commonly fixed rather than arising i.i.d. from a distribution. Consider the case where $Q^{\textup{{test}}}=\delta_{S}$ for some $S\in\mathcal{S}$. Using $k=1$ leads us to estimate the error using a single validation point, which is inconsistent due to observation noise at that point. Where the problem with the holdout estimator was bias, the problem with 1NN is variance. In section C.3.2 we prove proposition 4.4 and show 1NN has large error when applied to classification point prediction tasks.
Inconsistency of Nearest Neighbors When $k$ Is a Function of the Number of Validation Points. In fact, we can prove a more general result: that any nearest-neighbor test-risk estimator where the number of neighbors depends only on the number of validation points is inconsistent under infill asymptotics, regardless of type of dependence.
@@
Let $(k_{n})_{n=1}^{\infty}$ be any sequence of natural numbers. Define the sequence of estimators $\smash{\hat{R}_{N^{\textup{{val}}}}}$ to be the nearest neighbor risk estimators using $N^{\textup{{val}}}$ validation points and $k_{N^{\textup{{val}}}}$ neighbors. Then there exists a data-generating process satisfying infill asymptotics, a test set containing a single point, a predictive method $h$ resulting in an error function satisfying the Lipschitz assumption, and some $\!\epsilon,\delta>\!0$ such that with probability at least $1\!-\!\delta$, $\forall N^{\textup{{val}}}$, $|\smash{\hat{R}_{N^{\textup{{val}}}}}(h)-R_{Q^{\textup{{test}}}}(h)|\geq\epsilon$.
See section C.3.3 for a proof. There are two cases. (1) If the number of neighbors is bounded, the estimator can suffer from non-vanishing variance as in the 1NN case. Or (2) the number is unbounded, so there exists a sequence of validation sites that accumulates slowly around each test site leading to non-vanishing bias. Inconsistency of both 1NN and the holdout are corollaries of proposition 4.5; for 1NN, choose: $\forall n,k_{n}\!=\!1$. For the holdout, choose: $\forall n,k_{n}\!=\!n$.